# FitDiff: Robust monocular 3D facial shape and reflectance estimation using Diffusion Models

### v2

@@ 3.4 Sampling Procedure
DMs generate new samples by reversing the diffusion process, commencing from an initial random Gaussian noise. In parallel, the authors of  [83] introduced Denoising Diffusion Implicit Models (DDIMs) which are implicit probabilistic models [59]. They conduct a modified reverse diffusion process with fewer diffusion steps compared to those required during the vanilla DDPM sampling.
In our method, we adopt the DDIM sampling technique and integrate it into our trained architecture for the generation of facial avatars. We select the number of sampling steps to be $T=50$ for the DDIM sampling process. Aiming to generate accurate photorealistic avatars, we employ a novel guidance algorithm alongside the conditional input. Our approach is inspired by [19], in which the authors introduce a score corrector network conditioned on the diffusion step. In our implementation, we incorporate a face recognition network $\mathcal{C}$  [18] as a score corrector alongside with an off-the-shelf facial landmark detector $\mathcal{M}$ [12] and a perceptual loss [103]. The guidance method can be implemented for the vanilla DDPM [19] and the DDIM [83] sampling techniques. For the generation of intermediate images, we employ a differentiable renderer [75] with the modifications introduced in  [44, 45], under a single directional light. The pseudo-code and a detailed presentation of the proposed guidance method are presented in supplemental while the guidance formula is the following:
@@ 4.1 Dataset
Training such a diffusion model in a supervised manner requires a large dataset of labeled sets of facial images $\mathbf{I}$, facial textures $\mathbf{T}$, shape parameters $\mathbf{z}_{shp}$ and facial recognition embeddings $\mathbf{V}$. Although a captured dataset could be used, there are no large enough public datasets [106]. As a workaround, we curate 9000 2D facial images from the CelebA-HQ Dataset [37] $\mathbf{I}_{i},i=0\dots 9\cdot 10^{3}$, on which we performed the following steps to acquire the labeled dataset: A) We use a state-of-the-art face recognition model [18], to extract the identity latent embeddings $\mathbf{V}$, which captures the facial structure, with minimal interference from shading, age, and accessories. B) We train a state-of-the-art StyleGAN-based [40] facial reconstruction network [44] $\phi$, on public datasets of facial textures [63, 43], and use the LSFM 3DMM for the facial shape [8]. Following an iterative optimization [44], we fit our model to the CelebA-HQ dataset [37] and acquire pseudo-ground truth facial textures, 3DMM shape weights, and scene illumination parameters $\phi(\mathbf{I}_{i})\rightarrow\mathbf{A}_{D_{i}},\mathbf{A}_{S_{i}},\mathbf{N}_{i},\mathbf{z}_{shp_{i}},\mathbf{z}_{ill_{i}}$. In the end, we acquire a dataset of paired images, facial reflectance textures, facial shape, scene illumination and latent vectors: $\left\{\mathbf{I}_{i},\mathbf{A}_{D_{i}},\mathbf{A}_{S_{i}},\mathbf{N}_{i},\mathbf{z}_{shp_{i}},\mathbf{z}_{ill_{i}},\mathbf{V}_{i}\right\}$.
@@ 4.2 Unconditional Sampling
As mentioned in Sec. 3.3, FitDiff is trained following the classifier-free guidance (CFG) [34] training scheme. In this manner, our method can generate completely random facial identities without any prior input or supervision. We present the unconditional generated diffuse albedos $\mathbf{A}_{D}$, specular albedos $\mathbf{A}_{S}$, normals $\mathbf{N}$, facial shapes $\mathbf{S}$ and renderings in Fig. 5. This figure illustrates our method’s ability to create distinct shapes and textures. These assets hold significant potential for various applications, including enhancing existing datasets through augmentation and enrichment, as well as generating truly random identities for computer-based applications.
@@ 4.3 Qualitative comparisons
Input Ours  [64]  [44]  [55]  [45]  [27] Input Ours  [64]  [44]  [55]  [45]  [27]
Figure 6: Qualitative comparison between our method and other monocular-image facial reconstruction approaches [44, 55, 45, 27, 64]
We compare our method’s generated samples with other monocular-image face reconstruction methods [44, 55, 45, 27, 64] and present the generated samples in Fig. 6. The majority of these techniques rely on GAN-based methods and employ fitting optimization procedures that encompass lighting, camera pose, and expression parameters, whereas Relightify [64] is the only diffusion-based approach. Our method can capture finer details than most of GAN-based methods( [44, 55, 45, 27]) whereas it generates equally detailed avatars like Relightify.
@@ 4.4.2 Identity preservation experiment
Method ID Sim.
Label Only 0.43
CFG (w=2) 0.49
CFG (w=9) 0.45
Guidance 0.88
Table 2: Ablation study on the identity similarity [18] performance of our method, with and without identity guidance.
@@ 5.1 Sampling guidance
Firstly, we examine the importance of the proposed facial guidance during the reverse diffusion process. To do so, we randomly pick about 100 “in-the-wild" images, and we consider three scenarios: a) sampling without any guidance, b) sampling using the classifier-free guidance [34] while using guidance scales $w=\{2,9\}$, and c) our proposed method. The identity similarity scores are presented in Tab. 2, whereas visualizations of the generated examples are included in Fig. 9. The guidance algorithm demonstrates superior reconstruction performance compared to alternative methodologies.
@@
Figure 9: Ablation study: Given the input images (a), we present corresponding samples without the guidance algorithm (b), samples using CFG with w=2 (c) without the identity embedding(d) and finally our method (e).
5.2 Use of the conditioning mechanism
Another ablation study includes the necessity of our conditioning mechanism. We compare the texture information between the generated samples with and without our conditioning mechanism and examples of those are presented in Fig. 9d and 9e, respectively. These examples clearly show that finer details can be generated only when the corresponding identity embedding is used as input.
@@ 6 Discussion
FitDiff is a latent diffusion model, which concurrently carries out facial shape and texture generation as a combination of diffuse albedo, specular albedo, and shape normals. A close work to ours is Relightify [64], which only generates facial texture UV maps, modeled like ours. Given partially completed facial texture maps and acquired facial shape, this method can very efficiently reconstruct completed facial UV maps, through a diffusion-based sampling process. It treats the texture reconstruction problem as an in-painting approach, where off-the-shelve approaches are used to acquire the visible facial texture and geometry. Then, a multi-modal diffusion model is used to complete the non-visible parts of the facial texture maps, as well as reflectance. Relightify retains the visible input during inference, resulting in great performance in identity preservation (Sec. 4.4.2), which however introduces certain limitations: a) low-resolution images result in low-resolution textures, and b) wearable items (glasses, headbands) partially leak into the reflectance textures, as shown in Fig. 10.
On the contrary, FitDiff concurrently generates facial texture maps and geometry from scratch, and initializes the process from Gaussian noise. This results in a) robust shape and texture reconstruction even under occlusions and accessories, and b) always generating high-quality texture maps. Despite these advantages, our approach performs slightly worse than Relightify in facial texture benchmarks. We sincerely believe that the robustness benefits above make FitDiff an important alternative approach with multiple important use-cases. In fact, our model can actually be used in tandem with Relightify to provide more accurate shape priors to further boost its performance.
On the other hand, the very recent AlbedoGAN [72] is the first model that concurrently generates facial texture and shape. It provides FLAME [50] parameters alongside displacement maps for accurate facial shape reconstruction. While finetuning on the image input, it also produces a facial texture UV map often incorporating baked illumination and accessories. This consequently leads to reduced relightability in the generated avatars. It also replicates the limitations observed in Relightify, where wearable items are incorporated into the final facial texture(Fig. 10c). Finally, in comparison with FitMe [44], our experimentation results and qualitative assessments show that our generated texture maps achieve superior performance. We understand that this is due to the necessity in GAN-based fitting methods to meticulously initialize the GAN’s z or w embedding for back-propagation during inference, in order to mitigate optimization instabilities. The proposed method excels without the need for heuristic priors or regularization terms.
@@
Figure 10: Qualitative comparison with Relightify [64] and AlbedoGAN [72] with FitDiff. Relighitify is sensitive to occlusions such as glasses and headbands, while AlbedoGAN only produces one texture and essentially copies the input image, including glasses, teeth and shadows, while our method is robust to both and creates accurate and high-quality reflectance maps, including albedo, normals and specular.
@@ 7 Conclusion
In this paper, we introduced FitDiff, a diffusion-based 3D facial generative model, conditioned on identity embeddings. Acquiring such embeddings from a powerful pre-trained facial recognition method enables us to capture large variations in ethnicity, age, and gender from a single 2D facial image with no restriction on quality, pose, or illumination. Our method jointly generates facial shapes, facial reflectance maps, and scene illumination parameters. Through a series of experiments, FitDiff showcases state-of-the-art performance in preserving identity and reconstructing facial reflectance, either matching or surpassing established methods. Finally, it exhibits the capability to generate unconditional samples, further highlighting its versatility and effectiveness.
Acknowledgements:

### v3

@@ 3.4 Sampling Procedure
DMs generate new samples by reversing the diffusion process, commencing from an initial random Gaussian noise. In parallel, the authors of  [79] introduced Denoising Diffusion Implicit Models (DDIMs), which are implicit probabilistic models [56]. They conduct a modified reverse diffusion process with fewer diffusion steps than those required during the vanilla DDPM sampling.
In our method, we adopt the DDIM sampling technique and integrate it into our trained architecture to generate facial avatars. We select the number of sampling steps to be $T=50$ for the DDIM sampling process. Aiming to generate accurate photorealistic avatars, we employ a novel guidance algorithm alongside the conditional input. Our approach is inspired by [18], in which the authors introduce a score corrector network conditioned on the diffusion step. In our implementation, we incorporate a face recognition network $\mathcal{C}$  [17] as a score corrector alongside an off-the-shelf facial landmark detector $\mathcal{M}$ [11] and a perceptual loss [98]. The guidance method can be implemented for the vanilla DDPM [18] and the DDIM [79] sampling techniques. For the generation of intermediate images, we employ a differentiable renderer [71] with the modifications introduced in  [43, 44] under a single directional light. The pseudo-code and a detailed presentation of the proposed guidance method are presented in the supplemental while the guidance formula is the following:
@@ 4.1 Dataset
Training such a diffusion model in a supervised manner requires a large dataset of labeled sets of facial images $\mathbf{I}$, facial textures $\mathbf{T}$, shape parameters $\mathbf{z}_{shp}$ and facial recognition embeddings $\mathbf{V}$. Although a captured dataset could be used, there are no large enough public datasets [101]. As a workaround, we curate 9000 2D facial images from the CelebA-HQ Dataset [36] $\mathbf{I}_{i},i=0\dots 9\cdot 10^{3}$, on which we performed the following steps to acquire the labeled dataset: A) We use a state-of-the-art face recognition model [17], to extract the identity latent embeddings $\mathbf{V}$, which captures the facial structure with minimal interference from shading, age, and accessories. B) We train a state-of-the-art StyleGAN-based [39] facial reconstruction network [43] $\phi$, on public datasets of facial textures [60, 42], and use the LSFM 3DMM for the facial shape [7]. Following an iterative optimization [43], we fit our model to the CelebA-HQ dataset [36] and acquire pseudo-ground truth facial textures, 3DMM shape weights, and scene illumination parameters $\phi(\mathbf{I}_{i})\rightarrow\mathbf{A}_{D_{i}},\mathbf{A}_{S_{i}},\mathbf{N}_{i},\mathbf{z}_{shp_{i}},\mathbf{z}_{ill_{i}}$. After manually filtering out all failed cases, we finalize a dataset consisting of paired images, facial reflectance textures, facial shape, scene illumination and latent vectors: $\left\{\mathbf{I}_{i},\mathbf{A}_{D_{i}},\mathbf{A}_{S_{i}},\mathbf{N}_{i},\mathbf{z}_{shp_{i}},\mathbf{z}_{ill_{i}},\mathbf{V}_{i}\right\}$.
@@ 4.2 Unconditional Sampling
Following the classifier-free guidance (CFG) [33] training scheme, FitDiff can generate completely random facial identities without any prior input or supervision. We present the unconditional generated diffuse albedos $\mathbf{A}_{D}$, specular albedos $\mathbf{A}_{S}$, normals $\mathbf{N}$, facial shapes $\mathbf{S}$ and renderings in Fig. 4. This figure illustrates our method’s ability to create distinct shapes and textures.
@@ 4.3 Qualitative comparisons
Figure 5: Qualitative comparison between FitDiff and other monocular face reconstruction approaches [43, 52, 44, 26, 61].
We compare our method’s generated samples with other monocular-image face reconstruction methods [43, 52, 44, 26, 61] and present the generated samples in Fig. 5. Most of these techniques rely on GAN-based methods and employ fitting optimization procedures encompassing lighting, camera pose, and expression parameters, whereas Relightify [61] is the only diffusion-based approach. Our method can capture finer details than most GAN-based methods( [43, 52, 44, 26]), whereas it generates equally detailed avatars like Relightify.
@@ 5 Ablation Study
5.1 Conditioning embedding
The first ablation study focuses on the introduced conditioning embedding vector. Given the face recognition network [17], We examine 3 different types of input identity embeddings a) using only the last layer b) using only the first 3 layers and c) using all four of them, as proposed in FitDiff. We randomly pick about 100 “in-the-wild” images found across the internet, and we fit them using FitDiff. During sampling, we don’t use the guidance algorithm, as we want to let our network be entirely dependent on the identity embedding used. The first two rows of Table 2 show that the proposed condition vector surpasses the other.
@@ 5.2 Sampling guidance
Additionally, we examine the importance of the proposed facial guidance during the reverse diffusion process. Using the previously picked images, we consider three scenarios: a) sampling without any guidance, b) sampling using the classifier-free guidance [33] while using guidance scales $w=\{2,9\}$, and c) our proposed method. The identity similarity scores are presented in Tab. 2, whereas visualizations of the generated examples are included in Fig. 8. The guidance algorithm demonstrates superior reconstruction performance compared to alternative methodologies.
5.3 Use of the conditioning mechanism
Another ablation study includes the necessity of our conditioning mechanism. We compare the texture information between the generated samples with and without our conditioning mechanism, and examples of those are presented in Fig. 8d and 8e, respectively. These examples clearly show that finer details can be generated only when the corresponding identity embedding is used as input.
Method Using $n\in\{1,2,3\}$ layers Using $n$ layer FitDiff
ID Sim. 0.31 0.39 0.45
Method Label Only CFG (w=2) CFG (w=9) Guidance
ID Sim. 0.43 0.49 0.45 0.88
Table 2: Ablation study on the proposed conditioning input vector and the identity similarity performance of our method, using with and without identity guidance.
@@ 5.3 Use of the conditioning mechanism
Figure 8: Ablation study: From input images (a), we show results without the guidance algorithm (b), using CFG with w=2 (c), without identity embedding (d), and using our full method (e).
@@ 6 Discussion
(a) Input (b) Renderings (c) Input (d) Renderings
(a) Input (b) Rendering (c) Input (d) Rendering (c) Input (d) Rendering
Figure 9: Examples of FitDiff under extreme illumination (left) and extreme angles (right)
@@
Figure 10: Qualitative comparison between FitDiff, Relightify [61], and AlbedoGAN [68] highlights the strengths of our approach. While Relightify is sensitive to occlusions and AlbedoGAN generates a single texture by replicating the input image, FitDiff produces precise and high-quality facial reflectance maps.
(a) Input
(b) FitMe [43]
(c) FitDiff (Ours)
Figure 11: We applied our method to challenging images from FitMe [43], provided by the authors. Despite occlusions in the input facial images, our reconstructions remain unaffected.
FitDiff is a latent diffusion model which concurrently carries out facial shape and texture generation as a combination of diffuse albedo, specular albedo, and normals. A close work to ours is Relightify [61], which only generates facial texture UV maps modeled like ours. It treats the texture reconstruction problem as an in-painting approach by copying the visible part of the facial textures acquired from third-party off-the-shelve approaches. Then, a multi-modal diffusion model is used to complete the non-visible parts of the facial texture maps and reflectance. Relightify retains the visible input during inference, resulting in a great performance in identity preservation (Sec. 4.4.2), which, however, introduces certain limitations: a) low-resolution images result in low-resolution textures, and b) wearable items (glasses, headbands) partially leak into the reflectance textures, as shown in Fig. 11. On the contrary, FitDiff concurrently generates facial texture maps and geometry from scratch and initializes the process from Gaussian noise. This results in a) robust shape and texture reconstruction even under occlusions and accessories, and b) always generating high-quality texture maps. Despite these advantages, our approach performs slightly worse than Relightify in facial texture benchmarks. This is due to Relightify’s albedo-copying approach and the fact that FitDiff was trained in sorely synthetic data, contrary to the light-stage training data used in Relightify. We sincerely believe that the robustness benefits above make FitDiff an important alternative approach with multiple important use cases. In fact, our model can be used in tandem with Relightify to provide more accurate shape priors to further boost its performance.
On the other hand, the very recent AlbedoGAN [68] is the first model that concurrently generates facial texture and shape. It provides FLAME [49] parameters alongside displacement maps for accurate facial shape reconstruction. While finetuning on the image input, it also produces a facial texture UV map, often incorporating baked illumination and accessories. This consequently leads to reduced relightability in the generated avatars. It also replicates the limitations observed in Relightify, where wearable items are incorporated into the final facial texture(Fig. 11c).
Compared with FitMe [43], our experimentation results and qualitative assessments show that our generated texture maps achieve superior performance. This is due to the necessity in GAN-based fitting methods to meticulously initialize the GAN’s z or w embedding for back-propagation during inference to mitigate optimization instabilities. The proposed method excels without the need for heuristic priors or regularization terms.
@@ 7 Conclusion
In this paper, we introduced FitDiff, a diffusion-based 3D facial generative model conditioned on identity embeddings from a pre-trained facial recognition system. This approach captures diverse attributes like ethnicity, age, and gender from a single 2D image, with no restriction on quality, pose, or illumination. Our method jointly generates facial shapes, facial reflectance maps, and scene illumination parameters. Through a series of experiments, FitDiff showcases state-of-the-art performance in preserving identity and reconstructing facial reflectance, matching or surpassing established methods. Finally, it can generate unconditional samples and handles occlussions, further highlighting its versatility and effectiveness.
Acknowledgments: